# AI forecasting of higher-order wave modes of spinning binary black hole mergers

### v1

@@ Abstract
We present a physics-inspired transformer model that predicts the non-linear dynamics of higher-order wave modes emitted by quasi-circular, spinning, non-precessing binary black hole mergers. The model forecasts the waveform evolution from the pre-merger phase through to the ringdown, starting with an input time-series spanning $t\in[-5000\textrm{M},-100\textrm{M})$. The merger event, defined as the peak amplitude of waveforms that include the $\ell=|m|=2$ modes, occurs at $t=0\textrm{M}$. The transformer then generates predictions over the time range $t\in[-100\textrm{M},130\textrm{M}]$. Training, validation, and test sets were generated using the NRHybSur3dq8 model, considering a signal manifold defined by mass ratios $q\in[1,8]$; spin components $s^{z}_{\{1,2\}}\in[-0.8,0.8]$; modes up to $\ell\leq 4$, including the $(5,5)$ mode but excluding the $(4,0)$ and $(4,1)$ modes; and inclination angles $\theta\in[0,\pi]$. We trained the model on 14,440,761 waveforms, completing the training in 15 hours using 16 NVIDIA A100 GPUs in the Delta supercomputer. To evaluate the model, we utilized 4 H100 GPUs in the DeltaAI supercomputer to compute, within 7 hours, the overlap between ground truth and predicted waveforms on a test set of 840,000 waveforms. The mean and median overlaps achieved were 0.996 and 0.997, respectively. Additionally, we conducted interpretability studies to elucidate the waveform features that our transformer model utilizes to produce accurate predictions. The scientific software used for this work is released with this manuscript.
keywords:
AI , Transformers , Higher-order gravitational wave modes , Black Hole Mergers
journal: Physics Letters B
@@ 1 Introduction
Artificial Intelligence (AI) is being explored and applied in earnest to address computational grand challenges in gravitational wave astrophysics [1, 2, 3]. Contemporary AI applications cover the entire spectrum of gravitational wave data science, ranging from data quality [4, 5, 6, 7, 8], detection [9, 10, 11, 12, 13, 14, 15, 16, 17], parameter estimation [18, 19, 20, 21, 22, 23, 24], denoising [25, 26, 6, 27], detection [28, 29, 30, 31, 32] and forecasting [33, 34, 35] of sources with potential electromagnetic counterparts, among many others. Ref. [36] provides a comprehensive compilation of recent AI advances in gravitational wave astrophysics.
In this article we study whether AI can learn and predict the physics of higher-order wave modes emitted by the merger of quasi-circular, spinning, non-precessing binary black hole mergers. This work is inspired by previous studies demonstrating that AI can forecast the time-series signals of the $\ell=|m|=2$ mode of quasi-circular, non-spinning binary black hole mergers [37], and of quasi-circular, spinning, non-precessing binary black hole mergers [38]. In practice, these two studies [37, 38] demonstrated that AI can learn the key features that define the strongly non-linear, highly dynamical behaviour of binary black hole mergers, and then predict the time-series evolution of the pre-merger, merger and ringdown phases. This paper explores the use of transformers to learn the non-linear properties of higher-order wave modes and discusses methods for training these models on supercomputers with millions of waveforms.
To gain a basic understanding of the type of physics that physics-inspired AI models need to master in the context of higher-order wave modes, we present in Figure 1 a comparison between former studies and this work. The left column of Figure 1 displays samples of waveforms used in previous work  [37, 38], which have a very smooth evolution throughout inspiral, merger and ringdown. Although learning the physics of these waveforms poses challenges, these signals are relatively simple compared to the significantly more complex waveforms analyzed in this work, as illustrated in the right panel of Figure 1. In this case, the amplitude evolution is no longer monotonic up until merger, at $t=0\textrm{M}$, followed by a smooth decay. Rather, these waveforms show a complex structure due to the inclusion of higher-order wave modes, which also shift the location of the amplitude peak. For any set of parameters $(q,s_{1}^{z},s_{2}^{z})$, the inclination angle, $\theta$, modifies the physics of these waveforms significantly. For this reason, we densely sample the signal manifold with over 14 million waveforms to capture the dynamics of these higher-order wave modes. Our transformer model is designed to learn and forecast these intricate physical dynamics.
@@
Figure 1: Normalized amplitude $\left|h\right|/\left|h\right|_{\mathrm{max}}$ of gravitational waves around merger for different binary black hole configurations. Notice the impact of higher-order wave modes for inclination angles $\theta=0$ ( left column) and $\theta=\pi/4$ (right column).
As we demonstrate below, the algorithms we introduce in this work enable physics-inspired AI to learn and predict the nonlinear, highly dynamical evolution of higher-order wave modes with state-of-the-art accuracy. Key results of this work encompass:
  1. 1.
    A novel transformer model that is trained using the higher-order wave modes $\ell\leq 4$ and $(5,5)$, except for $(4,0)$ and $(4,1)$.
  2. 2.
    A new transformer architecture that takes as input data the two gravitational wave polarizations $\left(h_{+},h_{\times}\right)$, as given by $h(t)=h_{+}-ih_{\times}$, and forecasts time-series data for $\left(h_{+},h_{\times}\right)$ from the pre-merger phase through to ringdown.
  3. 3.
    A distributed training method that leverages 4 nodes, equivalent to 16 NVIDIA A100 GPUs, in the Delta supercomputer to train a transformer model with optimal forecasting performance within 15 hours using a training dataset of 14,440,761 million waveforms—the largest dataset used to date to train AI models for gravitational wave physics applications.
  4. 4.
    For inference, we used 4 NVIDIA H100 GPUs to quantify, within 7 hours, the performance of our transformer model using a test set of 840,000 waveforms. We found that the mean and median overlap between ground truth and predicted signals are 0.996 and 0.997, respectively, over the parameter space under consideration.
  5. 5.
    Open source scientific software to reproduce the key results of this article, available in GitHub [39].
@@
This paper is organized as follows. Section 2 describes the data used to train, validate and test our transformer model. It also provides a description of the physics-inspired design of our transformer model. We then discuss the training approach, and present inference results in Section 3. We summarize our key findings in Section 4. Future directions of work are outlined in Section 5. The numerical relativity waveforms used to train the NRHybSur3dq8 surrogate model use geometric units in which $G=c=1$, and M represents the total mass of the binary black-hole merger spacetime. Thus, the waveform datasets we use throughout this work follow the same convention, and we measure time in units of M, with $\textrm{M}=1\textrm{M}_{\odot}=4.93\times 10^{-6}\textrm{s}$.
@@ 2.1 Data
We produced three independent datasets of modeled waveforms using the numerical relativity (NR) surrogate model NRHybSur3dq8 [40]. These waveforms describe the gravitational emission from quasi-circular, spinning, non-precessing binary black hole mergers throughout the inspiral, merger, and ringdown phases. The datasets are constrained within the valid parameter space for this NR surrogate model, i.e., mass-ratios $q\leq 8$ and individual spins ($|s_{i}^{z}|\leq 0.8$ for $i=\{1,2\}$). The gravitational wave strain time-series, $h$, can be expressed as a sum of spin-weighted spherical harmonic modes, $h_{lm}$, on the 2-sphere [41]
@@
where ${}^{-2}Y_{lm}$ are the spin-weight-2 spherical harmonics, $\theta$ is the inclination angle between the orbital angular momentum of the binary and line of sight to the detector, and $\phi$ is the initial binary phase, which we set to zero. We include the higher-order wave modes $\ell\leq 4$ and $(\ell,m)=(5,5)$, except for (4,0) and (4,1). The waveforms cover the timespan $t=[-5000\textrm{M},130\textrm{M}]$.
Figure 2: Sampling of the signal manifold defined by mass-ratios $q\in[1,8]$, individual spins $s_{\{1,2\}}^{z}\in[-0.8,0.8]$, and inclination angle $\theta\in[0,\pi]$ to produce three independent datasets to train, validate and test our transformer models. The training set is produced using step sizes of $\{\delta q,\delta{s^{z}_{i}},\delta{\theta}\}=\{0.1,0.1,\pi/29\}$. The validation and test sets are generated by alternately sampling values between the training set values.
@@
The training dataset consists of 14,440,761 waveforms, generated by sampling the mass ratio $q\in[1,8]$ in steps of $\delta q=0.1$; individual spins $s_{i}^{z}\in[-0.8,0.8]$ in steps of $\delta s=0.1$; and inclination angle $\theta\in[0,\pi]$ in steps of $\delta\theta=\pi/29$.
@@
Each of these sets consists of 840,000 waveforms generated by sampling values spaced evenly between the training set values. Figure 2 illustrates the sampling strategy for generating the training, validation, and test sets, demonstrating that these datasets are independent and do not overlap.
Figure 3: Gravitational wave sequence and separation into encoder input, decoder input, and decoder target for training. Note that the decoder target complements the encoder input exactly, reflecting the transformer’s intended function of completing the signal. The decoder input sequence is shifted to the left by $\Delta t=1M$ with respect to the decoder target sequence and thus overlaps the encoder input by one timestep.
@@
We reduce the temporal resolution of the data by subsampling the original input sequences, selecting every other timestep. This approach reduces computational complexity. The modeled waveform is then split up such that the transformer’s encoder module receives the early inspiral sequence consisting of time steps $[-5000\textrm{M},-100\textrm{M}]$. The decoder module receives the inspiral, merger and ringdown segments made up of time steps $[-101\textrm{M},129\textrm{M}]$. The target waveform then consists of the gravitational wave segment $[-100\textrm{M},130\textrm{M}]$, i.e. the decoder target is one-time step ahead of the decoder input, as shown in Figure 3.
@@ 2.2 Architecture
Our transformer model is based on the original architecture introduced in Ref. [42], which leverages attention mechanisms to weigh different segments of the input sequence and improve performance in sequence prediction tasks. We have modified the model to incorporate specific properties of gravitational waves. Figure 4 schematically illustrates these modifications. Below, we describe the functionality of each component of the transformer as it processes the gravitational wave sequences. We represent the amplitudes of the input sequences for the encoder and decoder modules by $x_{p}$, where $p$ indicates the position within the sequence of length $n$. Note that the sequence length $n$ varies for encoder and decoder inputs, denoted as $n_{enc}$ and $n_{dec}$, respectively.
Figure 4: Schematic representation of transformer modules. The input time-series data for $(h_{+},h_{\times})$ is processed concurrently to generate corresponding predictions $(h_{+},h_{\times})$. This approach ensures the model maintains physical consistency in forecasting both time-series data
@@ 2.2.1 Embedding
The transformer architecture is not intrinsically sensitive to the order of tokens within the input sequence. However, for sequence prediction, information about token order is critical. This issue is addressed by introducing non-trainable positional encoding which embeds the input sequence into a higher-dimensional space. Through positional encoding, the input sequence $x_{p}$ is transformed into a sequence of $d_{embed}$ dimensional vectors each taking the form $X_{p}=(x_{p},$ $PE(p,1),$ $PE(p,2),$ ,...,… , $PE(p,d_{embed-1}))$. This transformation carries information about the original token’s position. We represent this transformed set of vectors as a matrix $X\in\mathbb{R}^{n\times d_{embed}}$. The embedding function $PE$, where $p$ is the position of any token within the original sequence and $i$ is the dimension within the positional encoding vector, takes the form:
@@
As demonstrated in Ref. [42], this type of sinusoidal positional encoding enables the model to interpolate and extrapolate positional encodings for unseen positions, providing better generalization than other embedding schemes. Building upon previous transformer models [38] for gravitational wave forecasting, this work facilitates the processing of complex-valued waveforms: Each input sequence is split into real and imaginary components, corresponding to plus, $h_{+}$, and cross, $h_{\times}$, polarizations, and independently embedded using the procedure outlined above. These embeddings are concatenated along the embedding dimension, effectively doubling the input’s dimensionality for subsequent processing by the encoder and decoder modules.
@@ 2.2.2 Encoder
Here, $Q^{j}\in\mathbb{R}^{n_{enc}\times d_{q}}$, $K^{j}\in\mathbb{R}^{n_{enc}\times d_{k}}$, and $V^{j}\in\mathbb{R}^{n_{enc}\times d_{v}}$ denote the query, key, and value matrices, respectively, defined as linear transformations of the input matrix $X$:
@@
with weight matrices $W_{Q}^{j}\in\mathbb{R}^{d_{embed}\times d_{q}}$, $W_{K}^{j}\in\mathbb{R}^{d_{embed}\times d_{k}}$, and $W_{V}^{j}\in\mathbb{R}^{d_{embed}\times d_{v}}$. In this work, $d_{k}$= $d_{q}$= $d_{v}$= $d_{embed}/\mathfrak{h}$.
$QK^{T}$ functions as a measure of similarity between keys and queries. The softmax activation operates on the rows of $QK^{T}$ and returns weights corresponding to values $V$. The output of each attention head is concatenated along the last dimension and then linearly transformed:
@@
with $W^{0}\in\mathbb{R}^{d_{embed}\times d_{embed}}$, enabling the model to integrate information across different representational subspaces. The resulting attention matrix has the same shape as the input $X$. After the attention mechanism, each position’s output undergoes processing by a position-wise feed-forward network, which applies two linear transformations with ReLU activations. Each sublayer (self-attention and feed-forward network) is enclosed by a residual connection, followed by layer normalization. This configuration aids in stabilizing the training process of deep networks.

### v2

# Sequence modeling of higher-order wave modes of quasi-circular, spinning, non-precessing binary black hole mergers

@@ Abstract
Higher-order gravitational wave modes from quasi-circular, spinning, non-precessing binary black hole mergers encode key information about these systems’ nonlinear dynamics. We model these waveforms using transformer architectures, targeting the evolution from late inspiral through ringdown. Our data is derived from the NRHybSur3dq8 surrogate model, which includes spherical harmonic modes up to $\ell\leq 4$ (excluding $(4,0)$, $(4,\pm 1)$ and including $(5,5)$ modes). These waveforms span mass ratios $q\leq 8$, spin components $s^{z}_{{1,2}}\in[-0.8,0.8]$, and inclination angles $\theta\in[0,\pi]$. The model processes input data over the time interval $t\in[-5000\textrm{M},-100\textrm{M})$ and generates predictions for the plus and cross polarizations, $(h_{+},h_{\times})$, over the interval $t\in[-100\textrm{M},130\textrm{M}]$. Utilizing 16 NVIDIA A100 GPUs on the Delta supercomputer, we trained the transformer model in 15 hours on over 14 million samples. The model’s performance was evaluated on a test dataset of 840,000 samples, achieving mean and median overlap scores of 0.996 and 0.997, respectively, relative to the surrogate-based ground truth signals. We further benchmark the model on numerical relativity waveforms from the SXS catalog, finding that it generalizes well to out-of-distribution systems, capable of reproducing the dynamics of systems with mass ratios up to $q=15$ and spin magnitudes up to 0.998, with a median overlap of 0.969 across 521 NR waveforms and up to 0.998 in face-on/off configurations. These results demonstrate that transformer-based models can capture the nonlinear dynamics of binary black hole mergers with high accuracy, even outside the surrogate training domain, enabling fast sequence modeling of higher-order wave modes.
@@ 1 Introduction
Artificial intelligence (AI) and machine learning (ML) are playing an increasingly central role in gravitational-wave (GW) astrophysics, offering powerful tools for tasks ranging from signal detection to population inference. Comprehensive reviews of these developments can be found in [1, 2]. These approaches have enabled the development of low-latency search pipelines [3], robust glitch classification [4, 5, 6], and accurate denoising methods [7]. Deep learning techniques have matched or exceeded traditional matched-filter performance in detecting binary black hole mergers [8, 9, 10, 11, 12, 13], and enabled efficient parameter estimation [14, 15] and population-level inference [16] orders of magnitude faster than classical methods.
[…]
@@
Figure 1: Normalized amplitude $\left|h\right|/\left|h\right|_{\mathrm{max}}$ of gravitational waves for different binary black hole configurations. Notice the impact of higher-order wave modes for inclination angles $\theta=0$ (left column) and $\theta=\pi/4$ (right column) in terms of amplitude modulations, and the time at which the waveform amplitude peaks.
[…]
In this work, we extend waveform forecasting to a significantly more complex regime: predicting both polarizations $(h_{+},h_{\times})$ across higher order harmonic modes, using only early inspiral information from quasi-circular, spinning, non-precessing binary black hole systems. This regime lies deep in the nonlinear sector of general relativity, where spin effects, mode coupling, and gravitational recoil are prominent, and analytical approximations break down. Higher-order modes are particularly important in systems with unequal masses, high spins, or edge-on orientations, where they enhance detection and parameter estimation accuracy [26, 27], but also introduce substantial modeling challenges [28]. Whereas previous studies [17, 19, 20, 22] focused on the quadrupole mode, we model both GW polarizations $(h_{+},h_{\times})$ and include all spherical harmonic modes up to $\ell=4$ (including $(\ell,m)=(5,5)$ and excluding $(4,0)$ and $(4,\pm 1)$, which are not reliably provided by the surrogate model used for training). Figure 1 illustrates the increased complexity introduced by these higher-order modes. The left panel shows representative waveforms from prior studies, where the amplitude evolves smoothly and is dominated by the quadrupole mode. In contrast, our target waveforms (right panel) show non-monotonic structure, strong modulation with inclination angle, and a shifted peak amplitude, demonstrating the richer dynamics that higher-order modes encode.
The key application of this transformer-based model is to serve as a fast surrogate for numerical relativity (NR) simulations, trained on physically realistic waveforms and designed to reflect key domain constraints such as polarization structure and sequence causality. Rather than functioning as a detection tool, it is designed for systematic modeling of higher-order modes, spin effects, and mode mixing across the BBH parameter space. This enables consistency tests of NR waveforms, waveform interpolation where NR is sparse or unavailable, and rapid hypothesis testing in gravitational theory. In doing so, it accelerates the development and validation of analytical waveform models and supports precision studies of strong-field gravity.
To implement this model, we introduce a set of architecture and training design choices that enable accurate and efficient sequence modeling of binary black hole waveforms:
1. Data generation and parameter space sampling We use the NRHybSur3dq8 surrogate model to generate gravitational waveforms with higher-order modes up to $\ell=4$ (including $(5,5)$ but excluding the $(4,0)$, $(4,\pm 1)$ modes, which the surrogate does not reliably support). The parameter space covers mass ratios $q\in[1,8]$, spins $s^{z}_{{1,2}}\in[-0.8,0.8]$, and inclination angles $\theta\in[0,\pi]$. To match this domain, we developed HPC-enabled sampling tools to densely cover the space, yielding over 14 million waveforms for training.
2. Transformer architecture and distributed training. To handle the vanishing $h_{\times}$ in edge-on mergers ($\theta=\pi/2$), we include conditionally activated processing layers that improve accuracy when $h_{\times}\approx 0$ due to the symmetry of the system. The model forecasts both GW polarizations, $(h_{+},h_{\times})$, from inspiral through ringdown as distinct channels, using a causal transformer that enforces smoothness and symmetry for physical consistency. Training using data parallelism across 16 NVIDIA A100 GPUs led to convergence within 15 hours on the Delta supercomputer.
3. Inference. Inference on 840,000 test waveforms was completed in under 5 hours using 1 NVIDIA V100 GPU on NCSA’s HAL system. The model achieved mean and median overlap scores of 0.996 and 0.997, respectively, when benchmarked against surrogate-model ground truth signals. When evaluated on a small set of numerical relativity waveforms beyond the training domain, the model achieves a mean overlap of 0.969 and a median of 0.902, which increased to a median of 0.998 and a minimum of 0.95 when NR waveforms were constrained to face-on binaries.
4. Exploratory interpretability studies. To better understand how the transformer achieves accurate predictions, we performed analyses examining the influence of key binary parameters—mass ratio, spin, and inclination angle—on model behavior. These exploratory studies provide insight into which waveform features the model finds predictive and how physical symmetries may be reflected in its learned representations. All code and analysis scripts are available at GitHub [29] to ensure transparency and reproducibility.
@@
This paper is structured as follows: Section 2 describes the data used to train, validate, and test the transformer model, and outlines the model design, training and inference methods. In Section 3, we present our findings. Future directions are discussed in Section 4. The numerical relativity waveforms employed in the NRHybSur3dq8 surrogate model use geometric units, where $G=c=1$ and M represents the total mass of the binary black-hole merger system. Consequently, all waveform datasets follow this convention, with time measured in units of M, where M is the total mass and $1\textrm{M}_{\odot}=4.93\times 10^{-6}$ seconds.
@@ 2 Methods
Here we describe the datasets used for this work, the approaches used to design and train our transformer model, and methods to test its predictive capabilities for sequence modeling.
@@ 2.1 Data
Three independent datasets of modeled waveforms were produced for training, validation, and testing of the transformer using the numerical relativity surrogate model NRHybSur3dq8 [21]. These waveforms describe the gravitational emission from quasi-circular, spinning, non-precessing binary black hole mergers, encompassing the inspiral, merger, and ringdown phases. The datasets are constrained within the valid parameter space for this NR surrogate model, i.e., mass ratios $q\leq 8$ and individual spins $|s_{\{1,2\}}^{z}|\leq 0.8$. The gravitational wave strain time-series $h(t,\theta,\phi)$ can be expressed as a sum of spin-weighted spherical harmonic modes, $h_{lm}$, on the 2-sphere [30]:
@@
where ${}^{-2}Y_{lm}$ are the spin-weight-2 spherical harmonics, $\theta$ is the inclination angle between the orbital angular momentum of the binary and the line of sight to the detector, and $\phi$ is the initial binary phase, which we set to zero. We include the higher-order wave modes $\ell\leq 4$ and $(\ell,m)=(5,5)$, except for $(4,0)$, $(4,-1)$ and $(4,1)$, since NRHybSur3dq8 only provides reliable extrapolation for these specific higher-order modes. The waveforms cover the time span $t=[-5000\,\textrm{M},130\,\textrm{M}]$.
Figure 2: Sampling of the signal manifold defined by the mass ratio $q\in[1,8]$, aligned spin components $s^{z}_{\{1,2\}}\in[-0.8,0.8]$, and inclination angle $\theta\in[0,\pi]$ to generate three independent datasets for training, validation, and testing of our transformer models. The training set uses a uniform grid with step sizes $\delta q=0.1$, $\delta s^{z}_{i}=0.02$, and $\delta\theta=\pi/29$. The validation and test sets are constructed by interleaving values between training grid points to avoid overlap and evaluate model generalization. For visual clarity, the figure shows a reduced interval of the full signal manifold.
@@
The training dataset consists of 14,440,761 waveforms, generated by sampling the mass ratio $q\in[1,8]$ in steps of $\delta q=0.1$; individual spins $s_{i}^{z}\in[-0.8,0.8]$ in steps of $\delta s^{z}=0.02$; and inclination angle $\theta\in[0,\pi]$ in steps of $\delta\theta=\pi/29$.
@@
Each of these datasets consists of 840,000 waveforms generated by sampling values spaced evenly between the training set values. Figure 2 illustrates the sampling strategy for generating the training, validation, and test sets, demonstrating that these datasets are independent and do not overlap.
Figure 3: Gravitational wave sequence and separation into encoder input, decoder input, and decoder target for training. Note that the decoder target complements the encoder input exactly, reflecting the transformer’s intended function of completing the signal. The decoder input sequence is shifted to the left by one timestep $\Delta t=1\,\textrm{M}$ with respect to the decoder target sequence and thus overlaps the encoder input by one timestep.
@@
We reduce the temporal resolution of the data by subsampling the original input sequences, selecting every other timestep. This approach reduces computational complexity. The modeled waveform is then split such that the transformer’s encoder module receives the early inspiral sequence, consisting of time steps $t=[-5000\,\textrm{M},-100\,\textrm{M}]$. The decoder module receives the inspiral, merger, and ringdown segments, made up of time steps $t=[-101\,\textrm{M},129\,\textrm{M}]$. The target waveform consists of the gravitational wave segment $t=[-100\,\textrm{M},130\,\textrm{M}]$, i.e., the decoder target is one timestep ahead of the decoder input, as shown in Figure 3.
@@ 2.2 Architecture
We extend the original architecture of the transformer model introduced in Ref. [31], which leverages attention mechanisms to weigh different segments of the input sequence and improve performance in sequence prediction tasks. Transformer-based architectures have recently shown strong performance in physical sciences, including stellar light curve analysis[32], gamma-ray burst afterglow modeling[33], and turbulent flow prediction [34]. Here, we have modified the model to incorporate properties of gravitational waves. These modifications include an embedding scheme that encodes both polarizations, $(h_{+},h_{\times})$, as well as an $h_{\times}$ mask that automatically activates additional layers for signals corresponding to edge-on mergers, which are difficult to resolve due to vanishing $h_{\times}$.
Figure 4 schematically illustrates these modifications. Below, we describe the functionality of each component of the transformer as it processes the gravitational wave sequences. We represent the amplitudes of the input sequences for the encoder and decoder modules by $x_{p}$, where $p$ indicates the position within the sequence of length $n$. Note that the sequence length $n$ varies for encoder and decoder inputs, denoted as $n_{\text{enc}}$ and $n_{\text{dec}}$, respectively.
Figure 4: Schematic representation of transformer modules. The input time-series data for $(h_{+},h_{\times})$ is processed concurrently to generate corresponding predictions $(h_{+},h_{\times})$. This approach ensures the model maintains physical consistency in forecasting both time-series data.
@@ 2.2.1 Embedding
The transformer architecture is not intrinsically sensitive to the order of tokens within the input sequence. However, for sequence prediction, information about token order is critical. This issue is addressed by introducing non-trainable positional encoding, which embeds the input sequence into a higher-dimensional space. Through positional encoding, the input sequence $x_{p}$ is transformed into a sequence of $d_{\text{embed}}$-dimensional vectors, each taking the form:
$X_{p}=(x_{p},PE(p,1),PE(p,2),\dots,PE(p,d_{\text{embed}-1})),$
where $PE(p,i)$ denotes the positional encoding at position $p$ and dimension $i$. This transformation carries information about the original token’s position, and we represent this transformed set of vectors as a matrix $X\in\mathbb{R}^{n\times d_{\text{embed}}}$. The embedding function $PE$, where $p$ is the position of any token within the original sequence and $i$ is the dimension within the positional encoding vector, takes the form:
@@
As demonstrated in Ref. [31], this type of sinusoidal positional encoding enables the model to interpolate and extrapolate positional encodings for unseen positions, providing better generalization than other embedding schemes. Building upon previous transformer models [20] for gravitational wave forecasting, this work facilitates the processing of complex-valued waveforms: Each input sequence is split into real and imaginary components, corresponding to plus, $h_{+}$, and cross, $h_{\times}$, polarizations, and independently embedded using the procedure outlined above. These embeddings are then concatenated along the embedding dimension, effectively doubling the input’s dimensionality for subsequent processing by the encoder and decoder modules.
@@ 2.2.2 Encoder
Here, $Q^{j}\in\mathbb{R}^{n_{\text{enc}}\times d_{q}}$, $K^{j}\in\mathbb{R}^{n_{\text{enc}}\times d_{k}}$, and $V^{j}\in\mathbb{R}^{n_{\text{enc}}\times d_{v}}$ denote the query, key, and value matrices, respectively, defined as linear transformations of the input matrix $X$:
@@
with weight matrices $W_{Q}^{j}\in\mathbb{R}^{d_{\text{embed}}\times d_{q}}$, $W_{K}^{j}\in\mathbb{R}^{d_{\text{embed}}\times d_{k}}$, and $W_{V}^{j}\in\mathbb{R}^{d_{\text{embed}}\times d_{v}}$. In this work, $d_{k}=d_{q}=d_{v}=d_{\text{embed}}/\mathfrak{h}$.
The matrix $QK^{T}$ functions as a measure of similarity between keys and queries. The softmax activation operates on the rows of $QK^{T}$ and returns weights corresponding to values $V$. The output of each attention head is concatenated along the last dimension and then linearly transformed:
@@
with $W^{0}\in\mathbb{R}^{d_{\text{embed}}\times d_{\text{embed}}}$, enabling the model to integrate information across different representational subspaces. The resulting attention matrix has the same shape as the input $X$. After the attention mechanism, each position’s output undergoes processing by a position-wise feed-forward network, which applies two linear transformations with ReLU activations. Each sublayer (self-attention and feed-forward network) is enclosed by a residual connection, followed by layer normalization. This configuration aids in stabilizing training.